OpenStudy (anonymous):

What is the height of the triangle if the area is one hundred eight square meters? eleven meters twelve meters sixty-nine meters seventy meters

OpenStudy (hoblos):

\[area = \frac{ base \times height }{ 2 }\] \[height = \frac{ 2 \times area }{ base }\]

OpenStudy (hoblos):

here area = 108 and base = 18 so what would the height be ?

OpenStudy (anonymous):